Red, Inc., Yellow Corp., and Blue Company each will pay a dividend of $3.05 next year. The growth rate in dividends for all three companies is 6 percent. The required return for each company's stock is 9.60 percent, 11.70 percent, and 14.40 percent, respectively.

    

Required:
(a) What is the stock price for Red. Inc., Company?
(Click to select)$31.77$9.15$19.55$50.83$84.72

    

(b) What is the stock price for Yellow Corp. Company?
(Click to select)$9.1526.07$50.83$53.51$17.23

      

(c) What is the stock price for Blue Company?
(Click to select)$36.31$50.83$21.18$9.15$14.95
Answer #1

before we start the calculations, let's recall the Gordan growth dividend discount model for calculating the share price. This model says:

Ae 9

Where, P = Price per share

D1 = Expected dividend next year = $ 3.05

Ke = Cost of equity or required return by the shareholder

g = perpetual growth rate = 6%

Part(a)

For Red Inc., Ke = 9.60%; Hence, P = 3.05 / (9.60% - 6%) = $ 84.72

Hence, the last option is the correct option.

---------------------------

Part (b)

For Yellow Corp., Ke = 11.70%; Hence, P = 3.05 / (11.70% - 6%) = $ 53.51

Hence the second last option is correct.

-----------------------

Part (c)

For Blue Company, Ke = 14.40%; Hence, P = 3.05 / (14.40% - 6%) = $ 36.31

Hence the first option is correct.